OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements S1, S2, and S3, and show that each of these statements is true. Show your work. Sn:   12 + 42 + 72 + . . . + (3n - 2)2 = n(6n^2-3n-1)/2

OpenStudy (anonymous):

Sn:   1^2 + 4^2 + 7^2 + . . . + (3n - 2)2 = n(6n^2-3n-1)/2

OpenStudy (anonymous):

@Agl202 any thoughts?

OpenStudy (anonymous):

first off can someone explain what the +...+

OpenStudy (zzr0ck3r):

So you want to show \[\sum_{i=1}^n(3i-2)^2=\frac{n(6n^2-3n-1)}{2}\]

OpenStudy (zzr0ck3r):

The \(+..+\) means it keeps going on in this way...

OpenStudy (anonymous):

where did you get the \[\sum_{i=1}^{n}\]

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

This is just what it means to add up a bunch of things.

OpenStudy (zzr0ck3r):

Do you know what induction is?

OpenStudy (anonymous):

ok and no please explain

OpenStudy (zzr0ck3r):

What class is this for, and how is it that you have a question that you seemingly know nothing about?

OpenStudy (anonymous):

pre calculus and because i have no teacher and am self teaching this

OpenStudy (zzr0ck3r):

Well the only way I know how to show this is through induction and I don't have time to teach you how to do all of that... Go read on induction and then you should be able to do this

OpenStudy (zzr0ck3r):

This seems like a little overboard for pre calc.

OpenStudy (anonymous):

could you just tell me what i need to do to solve it like what am i trying to get to

OpenStudy (zzr0ck3r):

Well \(\sum_{i=1}^n(3i-2)^2=1^2+4^2+7^2+10^2+13^2+...+(3n-2)^2\) We want to show \(\sum_{i=1}^n(3i-2)^2=\frac{n(6n^2-3n-1)}{2}\) for every natural number. To prove such things about the natural number, we use indiction. Induction says that if you can show something is true for \(n=1\) and you can show its true for \(n+1\) assuming it is true for \(n\) then it is true for all \(n\). For sure \(\sum_{i=1}^n(3i-2)^2=\frac{n(6n^2-3n-1)}{2}\) when \(n=1\) because \(\sum_{i=1}^1(3i-2)^2=1\) and \(\frac{1(6*1^2-3*1-1)}{2}=1\). Now we assume it's true for some \(n\) and want to show its true for \(n+1\). \(\sum_{i=1}^{n+1}(3i-2)^2= (3(n+1)-2)^2+\sum_{i=1}^n(3i-2)^2 = (3n+1)^2+\frac{n(6n^2-3n-1)}{2}=\\ \frac{(n+1)(6(n+1)^2-3(n+1)-1)}{2}\) and thus it is true for all \(n\).

OpenStudy (anonymous):

thank you

OpenStudy (zzr0ck3r):

np

OpenStudy (anonymous):

@zzr0ck3r one last thing i want to ask about this?

OpenStudy (zzr0ck3r):

yes?

OpenStudy (anonymous):

how does s1, s2, and s3 come into play

OpenStudy (zzr0ck3r):

This is a big question, google partial sums. http://www.mathsisfun.com/algebra/partial-sums.html

OpenStudy (anonymous):

@zzr0ck3r ok i understand it now i think so what does s1 stand for than?

OpenStudy (zzr0ck3r):

There is simply no way you read that... S1 is the first term S2 is the sum of the first 2 terms S3 is the sum of the first 3 terms . . . Sn is the sum of the first n terms

OpenStudy (anonymous):

so now im confused if sn is the statement i work with how do i write statment 1-3

OpenStudy (anonymous):

like the sum of what 2 terms for s2 for instance

OpenStudy (zzr0ck3r):

S2 = 1^2 + 4^2 S3 = 1^2+4^2+7^2 Sn = 1^2+4^2+7^2+10^2 + ... + (3n-2)^2

OpenStudy (anonymous):

thats what i thought

OpenStudy (zzr0ck3r):

note that 3(1) - 2 = 1 3(2) - 2 = 4 3(3)-2 = 7 3(4) - 2 = 10 . . .

OpenStudy (anonymous):

so since s2 is 1^2+4^2 would i plug nine into n for s2?

OpenStudy (anonymous):

im confused on what to do with the equation in connection to those terms

OpenStudy (zzr0ck3r):

ok so lets look at S3 1^2+4^2+7^2 = 1+16+49 = 66 now let \(n=3\) in \(\frac{n(6n^2-3n-1)}{n}\) and we get \(\frac{3(6*3^2-3*3-1)}{2}=66\)

OpenStudy (zzr0ck3r):

So this will always work for any n you choose.

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

so its finding n so it equals the sum of the three terms or how ever many we use!!!

OpenStudy (zzr0ck3r):

no, we dont find n, it works for all n

OpenStudy (zzr0ck3r):

but you would not want to show that it works for n = 23450987565746553738 because that would take your entire life, but since I proved it with a technique called induction we know its true for all n, including n = 23450987565746553738

OpenStudy (anonymous):

no i mean for s1 s2 and s3 its pluggin in a number for n that work for the sum of those terms used which intern says that all n works if it works for those

OpenStudy (anonymous):

your using sn to prove s1 s2 and s3 are true for this particular question

OpenStudy (zzr0ck3r):

no, that does not show it works for all n. The thing I did earlier shows that it works induction is basically this. if you can show that when something is true for n it must be true for n+1, and you can show its true for 1, then it must be true for 2 and then in turn it must be true for 3 and so on. So induction you do three things 1) show its true for n=1 2) assume its true for some arbitrary n 3) show that it must follow that it is true for n+1

OpenStudy (zzr0ck3r):

its like dominoes, if the one before you falls, then you are going to fall

OpenStudy (zzr0ck3r):

and if you start on the first domino, they will all fall

OpenStudy (anonymous):

thats what i mean though, is this question asking to prove sn to be true or for you to right the statements for term1 1 and than term 1+ term 2 and than term 1+ term 2 +term 3 showing what you must put into n to equal those sum of those terms

OpenStudy (anonymous):

im confused i think

OpenStudy (zzr0ck3r):

the question is asking you to show that for any \(n\) you choose the statement holds. Like I chose 3 and showed it held, and I could pick 4 and show it holds, and I could pick 27 and show it holds, but no matter how many cases I show it holds for does not prove that it holds for all cases. For that we need induction

OpenStudy (anonymous):

but what i dont get is for each statment i write how do i know what number to use for n

OpenStudy (zzr0ck3r):

I dont understand what you mean

OpenStudy (anonymous):

i have to write statement 1 statement 2 and statement 3. than show the statments are true so for instance s1 1^2= 1 so how do you know what to plug into for n so that (3n - 2)2 = n(6n^2-3n-1)/2 makes out to 1=1

OpenStudy (anonymous):

i mean thats easy because its obviously 1 that you plug into it

OpenStudy (anonymous):

but there other may not be

OpenStudy (zzr0ck3r):

this is not what you are asked to do and I don't know any better way to explain it. 1^2+4^2+7^2+10^2+......+(3n-2)^2 will always equal n(6n^2-3n-1)/2 no matter what n you pick.

OpenStudy (anonymous):

but i thought to write statements s1, s2 and s3 is just writing 1^2 1^2+4^2 1^2+4^2+7^2 and than prove that each of those is true

OpenStudy (anonymous):

and i had assumed to prove individually that those are true you had to plug into n to make the equation equal the sum of those

OpenStudy (zzr0ck3r):

yes and if you try and do that for every number in the world you are going to be here for an infinite amount of time

OpenStudy (anonymous):

ohh so i just write s1 s2 and s3 than state that using induction proves that those statments can be found in (3n - 2)2 = n(6n^2-3n-1)/2

OpenStudy (zzr0ck3r):

(3n - 2)2 = n(6n^2-3n-1)/2 only for n=1 but 1^2+4^2 = 2(6*2^2-3*2-1)/2 and 1^2+4^2+7^2 = 3(6*3^2-3*3-1)/2 and 1^2+4^2+7^2+10^2+14^2 +18^2 = 6(6*6^2-3*6-1)/2

OpenStudy (anonymous):

but thats what i mean how do you find that plugging in 2 into the n(6n^2-3n-1)/2 will equate to the sum of those two terms

OpenStudy (anonymous):

and the same for 3 instead of 2

OpenStudy (anonymous):

thats where im hung up on

OpenStudy (anonymous):

i mean what if you needed 67 instead of 2 how would you figure out you need 67

OpenStudy (zzr0ck3r):

\(\sum_{i=1}^3(3i-2)^2=1^2+4^2+7^2=\frac{3(6*3^2-3*3-1)}{2}\\ \sum_{i=1}^4(3i-2)^2=1^2+4^2+7^2+10^2=\frac{4(6*4^2-3*4-1)}{2}\\ \sum_{i=1}^5(3i-2)^2=1^2+4^2+7^2+10^2+14^2=\frac{5(6*5^2-3*5-1)}{2}\\ \sum_{i=1}^n(3i-2)^2=1^2+4^2+7^2+...+(3n-2)^2=\frac{n(6*n^2-3*n-1)}{2}\)

OpenStudy (zzr0ck3r):

you want to prove the last line

OpenStudy (zzr0ck3r):

ok good luck, I have to go. Just so you know. I did not learn anything about series (Sn) until calc 3, and I was not asked to prove things until 300 level math classes. I think you might be self studying things that you might not use for up to a year after you start calc, and thus are wasting your time.

OpenStudy (anonymous):

this is a pre calc class im studying for :(

OpenStudy (anonymous):

its part of my highschool curriculum